OpenStudy (anonymous):

The equation of a line is y=1/2x-2. What is the equation of the line that is parallel to the first line and passes through (4,-2)?

OpenStudy (anonymous):

***1/4

OpenStudy (anonymous):

how would you find the slope of the original line you're given?

OpenStudy (asnaseer):

are you saying the line apsses through (1/4,-2) or that the original line has an equation y=1/4x-2?

OpenStudy (asnaseer):

*passes

OpenStudy (anonymous):

A line in the form y=mx+b has a slope of m. so given that information, can you tell me the slope of that line, y=1/2x - 2?

OpenStudy (ash2326):

slope of 1st line =1/2 since 2nd line is parallel to first one then its slope is also 1/2 y=1/2x+c since it passes thjrough 4,-2 -2=1/2*4+c c=-4 y=1/2x-4